Titanium tetrachloride, TiCl4, is prepared by the reac- tion below.

3TiO2(s) + 4C(s) + 6Cl2(g) ¡

3 TiCl4(g) + 2 CO2(g) + 2 CO(g)

What is the maximum mass of TiCl4 that can be obtained from 35 g TiO2, 45 g Cl2, and 11 g C? 


1
Expert's answer
2021-09-23T02:39:16-0400

Solution:

The molar mass of TiO2 is 79.866 g/mol

The molar mass of Cl2 is 70.9060 g/mol

The molar mass of C is 12.0107 g/mol

The molar mass of TiCl4 is 189.679 g/mol


Balanced chemical equation:

3TiO2(s) + 4C(s) + 6Cl2(g) ⟶ 3TiCl4(g) + 2CO2(g) + 2CO(g)


Calculate moles of each reactant:

(35 g TiO2) × (1 mol TiO2 / 79.866 g TiO2) = 0.4382 mol TiO2

(11 g C) × (1 mol C / 12.0107 g C) = 0.9159 mol C

(45 g Cl2) × (1 mol Cl2 / 70.9060 g Cl2) = 0.6346 mol Cl2


Calculate stoichiometry coefficients:

TiO2: (0.4382 mol / 3 ) = 0.146

C: (0.9159 mol / 4) = 0.229

Cl2: (0.6346 mol / 6) = 0.106


As 0.106 is lowest, Cl2 is limiting reagent and product will form according to number of moles of Cl2.


According to stoichiometry:

6 mol of Cl2 produces 3 mol of TiCl4

Thus, 0.6346 mol of Cl2 produces:

(0.6346 mol Cl2) × (3 mol TiCl4 / 6 mol Cl2) = 0.3173 mol TiCl4


Calculate the mass of TiCl4:

(0.3173 mol TiCl4) × (189.679 g TiCl4 / 1 mol TiCl4) = 60.185 g TiCl4 = 60.2 g TiCl4


Answer: 60.2 grams of TiCl4